OpenStudy (anonymous):

Could someone explain to me how I can put this equation into my graphing calculator? I have tried everything, and I just don't understand it. y=-5secX I cant get it to graph. Any step by step explanation would help. Thank you!

OpenStudy (anonymous):

What graphing calculator do you have?

OpenStudy (anonymous):

I am using a TI-83 Plus

OpenStudy (anonymous):

I know how to add equation but since it SEC I am somehow not entering the correct thing

OpenStudy (anonymous):

Remember that sec(x)=1/cos(x)

OpenStudy (anonymous):

Try entering y=-5/cos(x) instead; it should produce the same graph.

OpenStudy (anonymous):

you are awesome! I actually see what I did wrong now. Thank you so much! Could you explain one more to me if you got time?

OpenStudy (anonymous):

I could try. :)

OpenStudy (anonymous):

y=cot(1/6x)

OpenStudy (anonymous):

cot(x)=1/tan(x)

OpenStudy (anonymous):

I got that far but what does it turn into? cot1/6x = 1/tan(6x) ?

OpenStudy (anonymous):

Is it (1/6)*x or 1/(6x)?

OpenStudy (anonymous):

y=cot(1/6x)

OpenStudy (anonymous):

cot(1/6x)=1/tan(1/6x)

OpenStudy (anonymous):

it gives me a wrong graph

OpenStudy (anonymous):

You have to be very careful using writing out what you mean, use a lot of parenthesis. :D What I wrote above can be interpreted two ways; 1/tan(1/(6x)) and 1/tan((1/6)*x)

OpenStudy (anonymous):

\[\frac{1}{\tan(\frac{1}{6x})}\] or \[\frac{1}{\tan(\frac{1}{6}x)}\]

OpenStudy (anonymous):

I tried both of them and neither looks like the correct graph on my homework. Its an online class and this is one of the questions I had wrong so I can see what the right graph is supposed to look like.

OpenStudy (anonymous):

tried to ask my prof. but she was no help at all!

OpenStudy (anonymous):

Maybe you got a strange scaling on?

OpenStudy (anonymous):

Usually with calculators you have to change the intervals.

OpenStudy (anonymous):

the pics on my calculator are nothing close to the image on the homework. let me try again tho..see if it changes anything

OpenStudy (anonymous):

Check if the boundaries are set proper; if this does not yield any enlightenment, I'm not sure I can help you further. :(

OpenStudy (anonymous):

..as I don't know what the problem could be then. :D

OpenStudy (anonymous):

No you did great! It helped. I changed the window a couple times and got it to look how it should. It wouldn't do it on its own I had to input the max and min myself! Thank you so much!! Hopefully I can do better on my test on Monday now! :)

OpenStudy (anonymous):

Glad I could help. Good luck with your test! :D

OpenStudy (anonymous):

Thank you! :)